# Google Tag Manager: Hidden Data Leaks and its Potential Violations under EU Data Protection Law

### 14 Dec 2023 v1

@@ 3.1.3. Artifacts and Responsible Disclosures
Artifacts associated to this work are available in an anonymous repository (supplementary-materials, nd). Additionally, two responsible disclosures are in progress, one with Google and another one with the Consentmanager CMP. An update will be added to this article in case of feedback from the companies.
@@ 3.2. Results
“He who says nothing agrees”: undefined consent variables are granted by default. In our experiments on consent mode and CMPs for Client-side GTM (see §3.1.1), we found that consent variables are in an undefined state when the CMP is loaded in the webpage. Though CMPs are expected to set default values to all such variables, we found out that some CMPs do not explicitly do that. Surprisingly, in this case, GTM considers all such undefined variables to be accepted by the end user, even though the end user has not interacted with the consent banner of the CMP yet. Among two CMPs tested (see §3.1.1), we detected this behavior for the Consentmanager CMP. This CMP sets a default value to only two consent variables – analytics_storage and ad_storage – leaving three GTM consent variables – security_storage , personalization_storage functionality_storage – and consent variables specific to this CMP – e.g., cmp_purpose_c56 which corresponds to the “Social Media” purpose – in undefined state. These extra variables are hence considered granted by GTM. As a result, all the Tags that depend on these four consent variables get executed even without user consent. We note that this default behavior of undefined variables cannot be changed and discuss its legal implications in the Potential violation 5 in Study II.
@@ 4.1.1. Personal data
It is “any information relating to an identified or identifiable natural person (’data subject’). An identifiable natural person is one who can be identified, directly or indirectly (GDPR, 2016, Art.4(11)). In order to determine whether a person is identifiable, account should be taken of all the means likely reasonably to be used by any actor to identify that person. Accordingly, if certain data, alone, is not personal data, it becomes personal data as regards to someone who reasonably has the means of enabling that data to be associated with a specific person (Sca, 2023, para. 46). GDPR Recital 30 asserts that online identifiers provided by their devices, such as IP addresses, can be associated to a person, thus making them identifiable. This identification does not require that all the information enabling that person to be identified should be in the hands of a single entity (Bre, 2016, para. 42, 43). Processing of personal data consists of ‘any operation(s) performed on personal data, such as collecting, sharing, using, making available, accessing, combining (GDPR, 2016, Article 4(2)). In practice, this means that almost any imaginable handling of personal data constitutes processing (29 Working Party, 2010; European Court of Justice, 2023).
@@ 4.1.2. Data Controller and Data Processor
According to European Data Protection Board (EDPB) opinion (29 Working Party, 2010), to determine whether an actor is a Data Controller, its factual roles and activities have to be evaluated in a specific situation. An actor is a Data Controller if it is responsible for determining the purposes, and the means of the processing of personal data (GDPR, 2016, Art. 4(7)).
“Determines” means having the “decision-making power” (European Data Protection Board, 2020; 29 Working Party, 2010; European Court of Justice, 2018a) or “independent control” (Information Commissioner’s Office, 2018) over the purposes and means of the processing (European Data Protection Board, 2020). The Court of Justice of the EU (CJEU), Data Protection Authorities (DPAs) and the EDPB describe that such control can be derived from: professional competence (legal or implicit) (29 Working Party, 2010); factual influence, based on the real circumstances surrounding the processing; image given to data subjects and their reasonable expectations on the basis of this visibility (29 Working Party, 2010); which actororganizes, coordinates and encourages” data processing (European Court of Justice, 2018a, para 70-71); and interpretation or independent judgement exercised to perform a professional service (Information Commissioner’s Office, 2018). “Purposes” refers to “why” data is processed. Purposes need to be explicit, specified and legitimate (GDPR, 2016, Art. 5(1)(b)). “Means” refers to the how the objectives of processing are achieved. The EDPB distinguishes between “essential” and “non-essential means” and provides examples thereof  (European Data Protection Board, 2020; 29 Working Party, 2010): Essential means are closely linked to the purpose and the scope of the processing and are inherently reserved to the controller. Examples of essential means are: determining the type of personal data processed; duration of processing, recipients of personal data, or categories of data subjects. Non-essential means may be delegated to the Data Processor, involving practical implementations, like hardware or software selection, security measures, or data storage/retrieval methods.
An actor is a Data Processor when it processes personal data on behalf of the data controller (GDPR, 2016, Art. 4(8)). An actor is a processor when: i) it is dependent on the controller’s instructions regarding processing activities  (29 Working Party, 2010)(European Data Protection Board, 2020), (Art. 28(3)(a)), Recital 81); and ii) complies with those instructions (European Data Protection Board, 2020).
@@ 4.1.3. Joint Controllership
Where two or more controllers jointly determine the purposes and means of processing, they shall be joint controllers (GDPR, 2016, Art. 26(1)). Joint participation can take the form of common or converging decisions on purposes and essential means (European Data Protection Board, 2020). Decisions can be considered as converging if they complement each other and are necessary for the processing to take place in such manner that they have a tangible impact on the determination of the purposes and means of the processing. An important criterion to identify converging decisions is whether the processing would not be possible without both parties’ participation in the sense that the processing by each party is inseparable, i.e., inextricably linked (European Data Protection Board, 2020). Joint controllership requires an agreement pursuant to GDPR Article 26. However, there is no requirement for both parties to share responsibility equally (European Court of Justice, 2018b, para 43). Similarly to Data Controllers, each of the both parties do not need to have access to personal data to be considered a controller(European Court of Justice, 2018b, para 38).
@@ 4.1.4. GDPR obligations for Data controllers and Data processors
  • Lawfulness: collect personal data only when it is legitimized with one legal basis (GDPR, 2016, Art. 5(1)(a)). In case this legal basis is consent, the consent must be prior to any data collection, freely given, specific, informed, unambiguous, readable, accessible, and revocable (GDPR, 2016, Art.4(11), Art. 7) (Santos et al., 2020);
  • Transparency: collect personal data when it informs the end user about the purposes, third party recipients, legal basis, among other information (GDPR, 2016, Art. 5(1)(a), 13, 14).
  • Security: ensure appropriate security of the personal data including protection against unauthorized or unlawful processing (GDPR, 2016, Art. 5(1)(f), 32).
  • Data Protection by Default: ensure that personal data is processed with the highest privacy protection so that by default personal data isn’t made accessible to an indefinite number of persons (GDPR, 2016, Art. 25(1)(2)).
  • Accountability (GDPR, 2016, Art.5(2)): the controller must be able to demonstrate compliance to each principle at any time;
@@ 4.2.1. Consent Management Platforms (CMPs)
Determination of Purposes. Through this scanning activity, the CMP extracts the purposes of the trackers from the Publisher’s website as shown in step (4a) of Figure 4. From reading the provided scanning report, the Publisher configures those purposes for each tag in the GTM interface as shown in step (5a) of Figure 4. As such, the Publisher determines the purposes for processing together with the CMP’s help. Notably, CMPs are argued to be recognized as Data Controllers, following the arguments of Santos et al. (Santos et al., 2021, Section 4.3). We posit that the final decision on the needed purposes for the Tag in question is a converging decision between Publishers and chosen CMPs (see §4.1.3), where both decisions to hire the CMP by the Publisher, and to identify the purposes by the CMP complement each other and are necessary for the processing to take place. Therefore, this decision (joint decision by the Publisher and the CMP) has a tangible impact on the determination of purposes of processing.
@@
Recommendation CMPs should either provide a comprehensive scanning service including all purposes both from stateless and stateful tracking technologies or avoid providing any scanner at all. Relatedly, we recommend GTM to provide to the CMP/Publisher the list of tags installed in the GTM Web and Server Containers so that CMPs will be able to reliably detect the Data Collectors and map the purposes of such tags.
@@
Mapping CMP purposes to GTM consent variables is not compliant. CMPs need to map GTM’s consent variables to the purposes shown in their consent banner. Notably, such mappings are not documented anywhere by CMPs nor by GTM. Furthermore, these mappings are subject to the legal issues identified in Potential Violation 3 and 4. When installing CookieBot and Consentmanager CMPs (claimed to be GTM compatible (Google, ndb)) on our website, we could infer the mappings of purposes (presented in Appendix Table 4 and 8). Herein we report several problems concerning such mappings: First, these CMPs use different names for purposes rather than the ones used by GTM. For example, the consent variable analytics_storage is called “Statistics” by Cookiebot, and “Measurement” by Consentmanager. The Publisher might not be aware that the purpose “Measurement” in the Consentmanager CMP banner corresponds to the analytics_storage consent variable in the GTM interface, and might have difficulty to configure tags that depend on this purpose. Second, certain CMPs merge multiple GTM purposes altogether, as displayed in Table 4. GTM consent variables functionality_storage and personalization_storage are merged into one category called “Preferences”. A Publisher does not know that the “Preferences” purpose in the Cookiebot refers to two different purposes in GTM which are personalization_storage and functionality_storage . Third, in our experiments we ascertained that Consentmanager only partially maps its purposes to GTM consent variables, as displayed in Table 8. Consentmanager does not map any of its own purposes to the ad_storage consent variable. If a user agrees to the marketing purpose shown in the consent banner, as there is no default consent variable that is associated with marketing purpose, a new variable (not included by default in GTM) is created within GTM. A Publisher might not be aware of this lack of mapping, neither of the need to configure such consent variable to a Tag. For instance, this means that a Publisher that uses the personalization_storage variable to configure a Tag, will see that this Tag will always run independently of user choices (see Technical finding 4).
@@ 4.2.2. GTM Provider
Determination of purposes. This actor is involved in the determination of purposes for two reasons. First, it defines five default consent variables in GTM “Consent Mode”: ad_storage , analytics_storage , security_storage , functionality_storage , and personalization_storage , presented in Table 3. We claim that these consent variables correspond to personal data processing purposes within the meaning of Article 5(1)(b) of the GDPR. A Publisher then needs to associate each Tag to one or more of these mentioned purposes, when purposes are not built-in. Second, we argue that GTM “organizes and coordinates” data processing (European Court of Justice, 2018a, paragraphs 70, 71), since it enables the orchestration of data collection and sharing through several actors - the Publisher, Tag Provider and the Data Collector. Hence, GTM Provider’s platform is the placeholder that coordinates the Tags for collecting data. GTM also encourages data processing by proposing the tags to Publishers.
@@
GTM purposes are not specific nor explicit. The GTM Provider, as a data controller, is obliged to provide specific, detailed and explicit purposes, according to the GDPR Purpose limitation principle (see Section 4.1.4). The description of the personalization_storage consent variable that says “storage related to personalization e.g. video recommendations” is not clear enough to understand to what final goal such personalization applies to - for example, if it is used for targeted advertising, it would require consent. These short and vague descriptions make it difficult for Publishers to conclude whether a purpose is exempted of consent (in case such purpose is functional or technical, see §4.1). Consequently, CMPs might run the risk to require consent for purposes exempted thereof. For example, Cookiebot requests consent for non-necessary purposes of “Preferences”, “Statistics” and “Marketing” in its consent banner. However, the “Preferences” purpose (see Table 9) seems to be necessary for a website to function, and hence, should be exempted of consent (ePD-09, 2009a, Art. 5(3))). As Cookiebot maps this purpose to the GTM consent variables personalization_storage and functionality_storage , according to their description in Table 3, it requests consent for this purpose.
Recommendation GTM should employ purpose-specific and explicit consent variables. Additionally, CMPs should not request consent for unnecessary purposes.
@@
Defaulting consent variables to “accepted” means that Tags run without consent. As described in the Technical Finding 4, the Web Container considers both “accepted” and “not defined” variables as accepted by the user, by default. We also discovered a case where the Consentmanager CMP allows some consent variables to be undefined. As such, the tags installed in the Web Container will run and collect user’s data without their due consent. This practice renders any processing of personal data potentially illegal (GDPR, 2016, Article 6(1)(a)). Notably, Consentmanager declares in its documentation that some consent mode functions are limited and not recommended for Publishers to use (Consentmanager, ndb). We argue that their disclaimer on the ”experimental mode” does not exempt them of accountability obligations (GDPR, 2016, Article 5(2)). Nevertheless, GTM still recommends to use this CMP without any warning, and it further states: “Discover featured CMP templates that deeply integrate with GTM’s consent configuration” in the interface. By doing so, GTM pushes Publishers to use a flawed CMP that might enable tags to run independently of consent, and thus enabling processing of personal data without legal basis.
@@ 4.2.3. Data Collector
Determination of Purposes The Data Collector, through the Terms and Conditions (T&Cs) or Privacy Policy it provides to the Publishers, decides upon the purposes and other conditions of processing, as shown in step (4b) in Figure 4. Examples for two specific case studies on two popular Data Collectors show how they communicate purposes to Publishers through T&Cs. In the documentation for Pinterest Tag, we found only one and very specific purpose of “online behavioural advertising” in the Advertising Guidelines(Pinterest, ndb). On its page, Pinterest says: “(…) information will be shared with third parties for online behavioural advertising”. In HotJar’s documentation, we found information about the purposes of the collected data in its data processing agreement (Data Processing Agreement, nd). In our interpretation, it pertains to the purposes of analytics and aggregated user behavior: “Hotjar allows its users to analyze and understand the behavioral patterns of their visitors (…). We reason there is a pluralistic control exercised by both the Data Collector and the Publisher: there is a common and complementary decision from both Publisher and Data Collector to process data, from the very moment in which the Publisher contractualizes with the Data Collector, derives the purposes from its T&Cs (Step 4b Fig. 4) and relies on the means of processing from the Data Collector. The processing would not be possible without the participation of both parties. This interdependence of the Publisher on the definition of purposes and means of the Data Collectors renders both actors joint controllers.

### 22 Dec 2023 v2

@@ 3.1.3. Artifacts and Responsible Disclosures
Artifacts associated to this work are available in an anonymous repository (supplementary-materials, nd). Additionally, one responsible disclosure is in progress with Google. An update will be added to this article in case of feedback from the company.
@@ 3.2. Results
“He who says nothing agrees”: undefined consent variables are granted by default. In our experiments on consent mode and CMPs for Client-side GTM (see §3.1.1), we found that consent variables are in an undefined state before the CMP is loaded in the webpage. Though CMPs are expected to set a default value to all such variables before the tags load, we found out that some CMPs do not do it fast enough. Surprisingly, in this case, GTM considers all such undefined variables to be accepted by the end user, even though the end user has not interacted with the consent banner of the CMP yet. Among two CMPs tested (see §3.1.1), we detected this behavior for the Consentmanager CMP, which correctly sets the analytics_storage and ad_storage before tag loading but define other CMP-specific variables such as cmp_purpose_c56 – which corresponds to the “Social Media” purpose – after tags already loaded. Since such variables are undefined when the tags are run, GTM considers them granted. As a result, all the Tags that depend on these variables get executed even without user consent. We note that this default behavior of undefined variables cannot be changed by the Publisher and discuss its legal implications in the Potential violation 5 in Study II.
@@ 4.1.1. Personal data
It is “any information relating to an identified or identifiable natural person (’data subject’). An identifiable natural person is one who can be identified, directly or indirectly (GDPR, 2016, Art. 4(11)). In order to determine whether a person is identifiable, account should be taken of all the means likely reasonably to be used by any actor to identify that person. Accordingly, if certain data, alone, is not personal data, it becomes personal data as regards to someone who reasonably has the means of enabling that data to be associated with a specific person (Sca, 2023, para. 46). GDPR Recital 30 asserts that online identifiers provided by their devices, such as IP addresses, can be associated to a person, thus making them identifiable. This identification does not require that all the information enabling that person to be identified should be in the hands of a single entity (Bre, 2016, para. 42, 43). Processing of personal data consists of ‘any operation(s) performed on personal data, such as collecting, sharing, using, making available, accessing, combining (GDPR, 2016, Art. 4(2)). In practice, this means that almost any imaginable handling of personal data constitutes processing (29 Working Party, 2010; European Court of Justice, 2023).
@@ 4.1.2. Data Controller and Data Processor
According to European Data Protection Board (EDPB) opinion (29 Working Party, 2010, para 20), to determine whether an actor is a Data Controller, its factual roles and activities have to be evaluated in a specific situation. An actor is a Data Controller if it is responsible for determining the purposes, and the means of the processing of personal data (GDPR, 2016, Art. 4(7)).
“Determines” means having the “decision-making power” (European Data Protection Board, 2020; 29 Working Party, 2010; European Court of Justice, 2018a) or “independent control” (Information Commissioner’s Office, 2018) over the purposes and means of the processing (European Data Protection Board, 2020). The Court of Justice of the EU (CJEU), Data Protection Authorities (DPAs) and the EDPB describe that such control can be derived from: professional competence (legal or implicit) (29 Working Party, 2010); factual influence, based on the real circumstances surrounding the processing; image given to data subjects and their reasonable expectations on the basis of this visibility (29 Working Party, 2010); which actororganizes, coordinates and encourages” data processing (European Court of Justice, 2018a, para. 70-71); and interpretation or independent judgement exercised to perform a professional service (Information Commissioner’s Office, 2018). “Purposes” refers to “why” data is processed. Purposes need to be explicit, specified and legitimate (GDPR, 2016, Art. 5(1)(b)). “Means” refers to the how the objectives of processing are achieved. The EDPB distinguishes between “essential” and “non-essential means” and provides examples thereof  (European Data Protection Board, 2020; 29 Working Party, 2010): Essential means are closely linked to the purpose and the scope of the processing and are inherently reserved to the controller. Examples of essential means are: determining the type of personal data processed; duration of processing, recipients of personal data, or categories of data subjects. Non-essential means may be delegated to the Data Processor, involving practical implementations, like hardware or software selection, security measures, or data storage/retrieval methods.
An actor is a Data Processor when it processes personal data on behalf of the data controller (GDPR, 2016, Art. 4(8)). An actor is a processor when: i) it is dependent on the controller’s instructions regarding processing activities  (29 Working Party, 2010)(European Data Protection Board, 2020),  (GDPR, 2016, Art. 28(3)(a), Recital 81 )); and ii) complies with those instructions (European Data Protection Board, 2020).
@@ 4.1.3. Joint Controllership
Where two or more controllers jointly determine the purposes and means of processing, they shall be joint controllers (GDPR, 2016, Art. 26(1)). Joint participation can take the form of common or converging decisions on purposes and essential means (European Data Protection Board, 2020) (paragraph 51). Decisions can be considered as converging if they complement each other and are necessary for the processing to take place in such manner that they have a tangible impact on the determination of the purposes and means of the processing. An important criterion to identify converging decisions is whether the processing would not be possible without both parties’ participation in the sense that the processing by each party is inseparable, i.e., inextricably linked (European Data Protection Board, 2020). Joint controllership requires an agreement pursuant to GDPR Article 26 (GDPR, 2016). However, there is no requirement for both parties to share responsibility equally (European Court of Justice, 2018b, para. 43). Similarly to Data Controllers, each of the both parties do not need to have access to personal data to be considered a controller(European Court of Justice, 2018b, para. 38).
@@ 4.1.4. GDPR obligations for Data controllers and Data processors
  • Lawfulness: collect personal data only when it is legitimized with one legal basis (GDPR, 2016, Art. 5(1)(a)). In case this legal basis is consent, the consent must be prior to any data collection, freely given, specific, informed, unambiguous, readable, accessible, and revocable (GDPR, 2016, Art. 4(11), 7) (Santos et al., 2020);
  • Transparency: collect personal data when it informs the end user about the purposes, third party recipients, legal basis, among other information (GDPR, 2016, Art. 5(1)(a), 13-14).
  • Security: ensure appropriate security of the personal data including protection against unauthorized or unlawful processing (GDPR, 2016, Art. 5(1)(f), 32);
  • Data Protection by Default: ensure that personal data is processed with the highest privacy protection so that by default personal data isn’t made accessible to an indefinite number of persons (GDPR, 2016, Art. 25(1)(2));
  • Accountability: the controller must be able to demonstrate compliance to each principle at any time (GDPR, 2016, Art. 5(2)).
@@ 4.2.1. Consent Management Platforms (CMPs)
Determination of Purposes. Through this scanning activity, the CMP extracts the purposes of the trackers from the Publisher’s website as shown in step (4a) of Figure 4. From reading the provided scanning report, the Publisher configures those purposes for each tag in the GTM interface as shown in step (5a) of Figure 4. As such, the Publisher determines the purposes for processing together with the CMP’s help. Notably, CMPs are argued to be recognized as Data Controllers, following the arguments of Santos et al. (Santos et al., 2021, §4.3). We posit that the final decision on the needed purposes for the Tag in question is a converging decision between Publishers and chosen CMPs (see §4.1.3), where both decisions to hire the CMP by the Publisher, and to identify the purposes by the CMP complement each other and are necessary for the processing to take place. Therefore, this decision (joint decision by the Publisher and the CMP) has a tangible impact on the determination of purposes of processing.
@@
Recommendation CMPs should provide a comprehensive scanning service including all purposes both from stateless and stateful tracking technologies. Relatedly, we recommend GTM to provide to the CMP/Publisher the list of tags installed in the GTM Web and Server Containers so that CMPs will be able to reliably detect the Data Collectors and map the purposes of such tags.
@@
Mapping between CMP purposes to GTM consent variables is not compliant. CMPs map the purposes shown in the CMP banner to consent variables in GTM. Notably, such mappings are not documented anywhereby Cookiebot nor GTM. Furthermore, these mappings are subject to the legal issues identified in Potential Violation 3 and 4. When installing CookieBot (claimed to be GTM compatible (Google, ndb)) on our website, we could infer the mappings of purposes 8). Herein we report several problems concerning such mappings: First, these CMPs use different names for purposes rather than the ones used by GTM. For example, the consent variable analytics_storage is called “Statistics” by Cookiebot, and “Measurement” by Consentmanager. The Publisher might not be aware that the purpose “Measurement” in the Consentmanager CMP banner corresponds to the analytics_storage consent variable in the GTM interface, and might have difficulty to configure tags that depend on this purpose. Second, certain CMPs merge multiple GTM purposes altogether, as displayed in Table 4. GTM consent variables functionality_storage and personalization_storage are merged into one category called “Preferences”. A Publisher does not know that the “Preferences” purpose in the Cookiebot refers to two different purposes in GTM which are personalization_storage and functionality_storage . Third, in our experiments we ascertained that Consentmanager only partially maps its purposes to GTM consent variables, as displayed in Table 8. Consentmanager does not map any of its own purposes to the ad_storage consent variable. If a user agrees to the marketing purpose shown in the consent banner, as there is no default consent variable that is associated with marketing purpose, a new variable (not included by default in GTM) is created within GTM. A Publisher might not be aware of this lack of mapping, neither of the need to configure such consent variable to a Tag. For instance, this means that a Publisher that uses the personalization_storage variable to configure a Tag, will see that this Tag will always run independently of user choices (see Technical finding 4).
@@ 4.2.2. GTM Provider
Determination of purposes. This actor is involved in the determination of purposes for two reasons. First, it defines five default consent variables in GTM “Consent Mode”: ad_storage , analytics_storage , security_storage , functionality_storage , and personalization_storage , presented in Table 3. We claim that these consent variables correspond to personal data processing purposes within the meaning of Article 5(1)(b) of the GDPR (GDPR, 2016). A Publisher then needs to associate each Tag to one or more of these mentioned purposes, when purposes are not built-in. Second, we argue that GTM “organizes and coordinates” data processing (European Court of Justice, 2018a, para. 70-71), since it enables the orchestration of data collection and sharing through several actors - the Publisher, Tag Provider and the Data Collector. Hence, GTM Provider’s platform is the placeholder that coordinates the Tags for collecting data. GTM also encourages data processing by proposing the tags to Publishers.
@@
GTM purposes are not specific nor explicit. The GTM Provider, as a data controller, is obliged to provide specific, detailed and explicit purposes, according to the GDPR Purpose limitation principle (see §4.1.4). The description of the personalization_storage consent variable that says “storage related to personalization e.g. video recommendations” is not clear enough to understand to what final goal such personalization applies to - for example, if it is used for targeted advertising, it would require consent. These short and vague descriptions make it difficult for Publishers to conclude whether a purpose is exempted of consent (in case such purpose is functional or technical, see §4.1). Consequently, CMPs might run the risk to require consent for purposes exempted thereof. For example, Cookiebot requests consent for non-necessary purposes of “Preferences”, “Statistics” and “Marketing” in its consent banner. However, the “Preferences” purpose (see Table 9) seems to be necessary for a website to function, and hence, should be exempted of consent (ePD-09, 2009a, Art. 5(3))). As Cookiebot maps this purpose to the GTM consent variables personalization_storage and functionality_storage , according to their description in Table 3, it requests consent for this purpose. Among the two CMPs we have tested, this risk only affects the Cookiebot CMP.
Recommendation GTM should employ purpose-specific and explicit consent variables. Additionally, Cookiebot should not request consent for unnecessary purposes.
@@
Defaulting consent variables to “accepted” means that Tags run without consent. As described in the Technical Finding 4, the Web Container considers both “accepted” and “not defined” variables as accepted by the user, by default. We also discovered a case where the Consentmanager CMP allows some consent variables to be undefined. As such, the tags installed in the Web Container will run and collect user’s data without their due consent. This practice renders any processing of personal data potentially illegal (GDPR, 2016, Art. 6(1)(a)). Notably, Consentmanager declares in its documentation that some consent mode functions are limited and not recommended for Publishers to use (Consentmanager, ndb). We argue that such disclaimer does not exempt CMPs of accountability obligations (GDPR, 2016, Art. 5(2)). Nevertheless, GTM still recommends to use this CMP without any warning, and it further states: “Discover featured CMP templates that deeply integrate with GTM’s consent configuration” in the interface. By doing so, GTM pushes Publishers to use a CMP that might enable tags to run independently of consent, and thus enabling processing of personal data without legal basis.
@@ 4.2.3. Data Collector
Determination of Purposes The Data Collector, through the Terms and Conditions (T&Cs) or Privacy Policy it provides to the Publishers, decides upon the purposes and other conditions of processing, as shown in step (4b) in Figure 4. Examples for two specific case studies on two popular Data Collectors show how they communicate purposes to Publishers through T&Cs. In the documentation for Pinterest Tag, we found only one and very specific purpose of “online behavioural advertising” in the Advertising Guidelines (Pinterest, ndb). On its page, Pinterest says: “(…) information will be shared with third parties for online behavioural advertising”. In HotJar’s documentation, we found information about the purposes of the collected data in its data processing agreement (Data Processing Agreement, nd). In our interpretation, it pertains to the purposes of analytics and aggregated user behavior: “Hotjar allows its users to analyze and understand the behavioral patterns of their visitors (…)”. We reason there is a pluralistic control exercised by both the Data Collector and the Publisher: there is a common and complementary decision from both Publisher and Data Collector to process data, from the very moment in which the Publisher contractualizes with the Data Collector, derives the purposes from its T&Cs (Step 4b Fig. 4) and relies on the means of processing from the Data Collector. The processing would not be possible without the participation of both parties. This interdependence of the Publisher on the definition of purposes and means of the Data Collectors renders both actors joint controllers.